Are you racialy G rachel dolezal 20 of the most sp Wetch Online ScandeE Watch the O lon capaauamed net/enc/48/f6cd3b318144164cSadfd7e2a6f2e5col6249947a4 Ashley Forde (Student) Physics 2-Spring 2018 Course Contents ...Chapter 20 AC Current A plot of current in mA as a function of time in seconds is shown above. What is the angular frequency (e) of the current? 0.45 rads Submit Answer Incorrect Tries 2/10 Previous Tries Sibmit Anwer Incorrect Tries 1/10 Previous Tries 3 5 caps lock
Answer #1

(a) in 11 sec it completes 2.5 cycle

So f = 2.5/11

\omega = 2πf = 5π/11 = 1.43rad/sec

(B) irms = imax./√2 = 7.5/√2 = 5.303 mAmp =0.0053Amp